OpenStudy (anonymous):

Is 7/9 closer to 1/2 or 1?

OpenStudy (anonymous):

7/9 is closer to 1 because first you can't simple 7/9 to 1/2 and 7 is close to 9 and 9/9 is a whole number so that is why I think 7/9 is closer to 1 than 1/2. but i can be wrong, so sorry if i am wrong

OpenStudy (anonymous):

@HelpIsuckatmath39

OpenStudy (anonymous):

 you can see now or 1/2=0.50 7/9=0.78 1=1.00 1-0.78=0.12 0.78-0.50=0.28 you can see now.